Reports 1-1 of 1 Clear search Modify search
satoru.ikeda - 11:46 Friday 12 March 2021 (16248) Print this report
Comment to Monitoring of IP by ITMY venting (16091)

I have confirmed the problem reported in KLog#16115, where H2 was attached.

As you can see in the photo, it was attached to the limit switch on the R side of H2.
The other H1 and H3 were not attached to the limit switch.

However, H1, H2, and H3 are all moving in the same direction, and we believe that they are rotating in the YAW direction.

We will discuss the release of the limit switches at a later date.

Images attached to this comment
Non-image files attached to this comment
Search Help